# Synchronized Video Storytelling: Generating Video Narrations with Structured Storyline

### v1

@@ Abstract
Video storytelling is engaging multimedia content that utilizes video and its accompanying narration to attract the audience, where a key challenge is creating narrations for recorded visual scenes. Previous studies on dense video captioning and video story generation have made some progress. However, in practical applications, we typically require synchronized narrations for ongoing visual scenes. In this work, we introduce a new task of Synchronized Video Storytelling, which aims to generate synchronous and informative narrations for videos. These narrations, associated with each video clip, should relate to the visual content, integrate relevant knowledge, and have an appropriate word count corresponding to the clip’s duration. Specifically, a structured storyline is beneficial to guide the generation process, ensuring coherence and integrity. To support the exploration of this task, we introduce a new benchmark dataset E-SyncVidStory with rich annotations. Since existing Multimodal LLMs are not effective in addressing this task in one-shot or few-shot settings, we propose a framework named VideoNarrator that can generate a storyline for input videos and simultaneously generate narrations with the guidance of the generated or predefined storyline. We further introduce a set of evaluation metrics to thoroughly assess the generation. Both automatic and human evaluations validate the effectiveness of our approach. Our dataset, codes, and evaluations will be released.
@@
Dingyi Yang1, Chunru Zhan2, Ziheng Wang1, Biao Wang2, Tiezheng Ge2, Bo Zheng2, Qin Jin1* 1School of Information, Renmin University of China 2Alibaba Group {yangdingyi,zihengwang,qjin}@ruc.edu.cn {zhanchunru.zcr,eric.wb,tiezheng.gtz,bozheng}@alibaba-inc.com
@@ 1 Introduction
Figure 1: An example of our annotated synchronized video storytelling, with English translations provided for easy reading. Sequential visual scenes are narrated, following a storyline expressed by categorized script labels. Each narration should incorporate appropriate relevant knowledge (as underlined in the figure) and have a character count that fits the duration of each clip.
@@
In this paper, we introduce the task of Synchronized Video Storytelling, which aims to generate synchronized narrations for videos. As illustrated in Figure 1, the sequential visual scenes in the video are accompanied by corresponding narrations. Each narration has a suitable word sequence length and incorporates relevant knowledge to effectively engage the audience. They can be directly utilized as voiceovers for video recordings. This distinguishes our task from existing ones, as illustrated in Figure 2. Furthermore, our generation follows a structured storyline, similar to the human writing process, beginning with a brief outline and then drafting the complete story Yang et al. (2022b). This proposed task is complex and cannot be effectively solved by traditional frameworks or evaluated by existing benchmarks. A possible approach is to utilize powerful Multi-modal Large Language Models (MLLMs) Han et al. (2023); Sun et al. (2023) for zero-shot or few-shot generation. However, most MLLMs face challenges with time alignment and sequential video comprehension Huang et al. (2023). Moreover, generating narrations for longer videos becomes difficult for models due to the requirement of excessively long visual tokens. Another approach is to generate captions for video shots and use them as input for the LLMs Bhattacharya et al. (2023); Lin et al. (2023b). However, this relies heavily on the performance of the captioning model. Moreover, current models are ineffective in generating coherent, appealing, and length-appropriate narrations. Additional training on specific instruction datasets is necessary.
To support the exploration of the proposed task, we create a benchmark dataset named E-SyncVidStory (E-commerce Synchronized Video Storytelling). Given the rapid growth of e-commerce platforms, many sellers have product information and visual materials, but may lack the time and resources to convert these into engaging advertisement videos. Therefore, they seek an automatic and efficient solution. Our dataset can offer both substantial research and practical value in this area. E-SyncVidStory contains 6.0K videos with 41.3K video clips in total. Each video is segmented as sequential visual scenes with synchronized Chinese narrations. As the primary objective of Ad stories is to engage with the audience and showcase products Bhattacharya et al. (2023), they might not always present a sequence of events as a classic story. Still, like classic ones, advertising stories should be coherent and follow a logical storyline, which is annotated in our dataset [^1] The storyline is expressed by categorized script labels. These labels are divided into 12 types, with their definitions provided in Appendix A.4.. All annotations are obtained through automatic preprocessing, GPT-based refinement, and manual checks.
@@ 3.1 Task Definition
Given a video $V$ composed of sequential video clips $\{v_{1},v_{2},...,v_{n}\}$ and the related knowledge composed of several knowledge items $\mathcal{K}=\{k_{1},k_{2},...,k_{m}\}$, synchronized video storytelling aims to incorporate the knowledge to generate a narration $s_{i}$ for each video clip $v_{i}$, and all narrations form a coherent story $\mathbf{S}=\{s_{1},s_{2},...,s_{n}\}$ for the video. The word count  [^2] If the time length of video clip $v_{i}$ is $t$ seconds, considering the normal speaking speed of 5 words per second, the appropriate word length range is $[5t-3,5t+3]$ with an error margin of 3 words in either direction. of each $s_{i}$ should be within the appropriate range $r_{i}$. To enhance the coherence of the overall story, the generation is guided by a predefined or automatically predicted storyline, which is expressed using a sequence of script labels $\{l_{1},l_{2},...,l_{n}\}$. Each script label $l_{i}$, belonging to one of the 12 predefined types, summarizes the key point of the corresponding narration. More details about the script label are presented in the Appendix A.4.
@@ 4.1 Visual Embedding
For long videos consisting of multiple video clips, when generating a narration for the current clip $v_{i}$, we only need a more concise message of previous video clips. Inspired by Song et al. (2023), we perform memory consolidation by merging the most similar visual tokens. Specifically, we maintain fixed-length tokens as visual memory. We iteratively repeat the merging process until the memory tokens of previous video clips reach the fixed length.
@@
4.2 Prompting
The input prompt for VideoNarrator is displayed in Appendix Figure 6. For each video clip, the controllable signal is provided just before the narration, which can ensure controllable accuracy. In our instruction, we concisely explain the task requirements. As shown in Figure 6, these red words can make it easier for the LLM to target relevant information. In the underlined instruction, we emphasize that the model should primarily use the information provided in the prompt and avoid using unrelated knowledge stored in the LLM.
@@ 5.1 Experiment Settings
In our experiments, we use pre-trained models as visual feature extractors as described in the above Section. For the LLM model, we employ the Baichuan-7B model Yang et al. (2023), known for its excellent performance in Chinese-related tasks. We construct our dataset as illustrated in Figure 6, resulting in 41k instruction training samples. Our proposed framework is trained for 15 epochs, using a learning rate of $1e^{-4}$ and a batch size of 32. The LoRA parameters are set to $r=64$ and $alpha=16$. We apply 90% of videos as the training set and the left as the testing set. The training of our 7B model took around $48$ hours on 4 A6000 GPUs.
@@
We compare our model with the following baselines. (1) Multi-model Pipieline. Specifically, we apply LLaVA Liu et al. (2023) to get textual summaries for video clips and prompting GPT-3.5 for zero-shot and few-shot generation. (2) End2end MLLMs, including zero-shot performance of Video-ChatGPT Maaz et al. (2023a), Video-LLaVA Lin et al. (2023a), and VTimeLLM Huang et al. (2023) with 7B LLMs. These models take the video as inputs and generate narrations in an end2end manner. (3) Fine-tuned End2end MLLMs. Specifically, we finetune the VTimeLLM model, which is effective for dense video comprehension and supports Chinese generation. More details about the baselines can be found in our Appendix A.1.
@@
Figure 7: Examples of stories generated by our proposed VideoNarrator.
@@ 5.2 Automatic Evaluation Results
With Generated Storyline. As shown in Table 2, our model outperforms the baseline models, even when fine-tuned on our dataset. For detailed evaluations shown in Table 3, existing MLLMs demonstrate their ability to generate visually relevant and informative stories. However, they tend to focus on a limited amount of prior knowledge and may repeat it, resulting in a lower $\textrm{Info}_{\textrm{Diverse}}$ score. Furthermore, their fluency and ability to control the length of word sequences are poor. Based on the ablation studies, we find that by fine-tuning the LLM with our instruction data, our model can achieve the ability to control the word sequence length and comprehend video features, showing an increase in controllable accuracy and visual relevance. Applying visual compression and visual memory consolidation can improve the model’s ability to extract relevant knowledge and retain story coherence, resulting in better performance in knowledge relevance and fluency. Explicit storylines and relative positional embedding can also improve the performance in knowledge relevance and fluency.
@@
To validate the effectiveness of our proposed VideoNarrator, we conduct experiments on the Video Storytelling dataset Li et al. (2019), which covers a more general domain. We make minor adjustments to the dataset to fit the task setting of Synchronized Video Storytelling. Specifically, we apply the text length of the ground truth as the length requirement, and use the video topic as its relevant knowledge. The training prompt is similar to that presented in Figure 6, while the system role is changed to “You are an AI assistant that can understand videos”, the product information is replaced with the event topic, and the requirement for script labels has been removed. Additionally, the LLM model is replaced by LLaMA-2 Touvron et al. (2023) for English generation. As shown in Table 5, our model surpasses the state-of-the-art results. This showcases its effectiveness in synchronized video storytelling across a broader domain.
@@ 6 Conclusion
This paper introduces the new task of synchronized video storytelling, which aims to generate synchronous narrations for sequential video scenes. These narrations are guided by a structured storyline and can incorporate relevant knowledge, resulting in a coherent and informative story. The new task is more practical than existing tasks of dense video captioning and visual storytelling because the generated stories can be directly combined with the visual scenes, creating an engaging video-format story. We collect a benchmark dataset called E-SyncVidStory and introduce an effective framework named VideoNarrator. Both automatic and human evaluations verify the effectiveness of our proposed framework.
@@
After refining the data using the powerful GPT-4, we only need to conduct further checking of the annotations. We have hired 5 workers to review and correct any remaining errors. It took approximately 2 minutes for each video, and the workers were paid 1 RMB per video, which is a reasonable payment for local people.
@@ A.2 Evaluation Metric for Knowledge Relevance
where $W(s_{i})$ represents all the words in sentence $s_{i}$. $f_{s_{i}},f_{k}$, and $f_{w}$ refer to the normalized embeddings of sentence $s_{i}$, knowledge point $k$, and segmented word $w$, respectively.

### v2

@@ Abstract
Video storytelling is engaging multimedia content that utilizes video and its accompanying narration to attract the audience, where a key challenge is creating narrations for recorded visual scenes. Previous studies on dense video captioning and video story generation have made some progress. However, in practical applications, we typically require synchronized narrations for ongoing visual scenes. In this work, we introduce a new task of Synchronized Video Storytelling, which aims to generate synchronous and informative narrations for videos. These narrations, associated with each video clip, should relate to the visual content, integrate relevant knowledge, and have an appropriate word count corresponding to the clip’s duration. Specifically, a structured storyline is beneficial to guide the generation process, ensuring coherence and integrity. To support the exploration of this task, we introduce a new benchmark dataset E-SyncVidStory with rich annotations. Since existing Multimodal LLMs are not effective in addressing this task in one-shot or few-shot settings, we propose a framework named VideoNarrator that can generate a storyline for input videos and simultaneously generate narrations with the guidance of the generated or predefined storyline. We further introduce a set of evaluation metrics to thoroughly assess the generation. Both automatic and human evaluations validate the effectiveness of our approach. Our dataset, codes, and evaluations are released at https://github.com/alibaba/alimama-video-narrator.
@@
Dingyi Yang1, Chunru Zhan2, Ziheng Wang1, Biao Wang2, Tiezheng Ge2, Bo Zheng2, Qin Jin1* 1Renmin University of China 2Alibaba Group {yangdingyi,zihengwang,qjin}@ruc.edu.cn {zhanchunru.zcr,eric.wb,tiezheng.gtz,bozheng}@alibaba-inc.com
@@ 1 Introduction
Figure 1: An example of our annotated synchronized video storytelling, with English translations provided for easy reading. Sequential visual scenes are narrated, following a storyline expressed by categorized script labels. Each narration should incorporate appropriate relevant knowledge (as underlined in the figure) and have a word count that fits the duration of each clip.
@@
In this paper, we introduce the task of Synchronized Video Storytelling, which aims to generate synchronized narrations for videos. As illustrated in Figure 1, the sequential visual scenes in the video are accompanied by corresponding narrations. Each narration has a suitable word sequence length and incorporates relevant knowledge to effectively engage the audience. They can be directly utilized as voiceovers for video recordings. This distinguishes our task from existing ones, as illustrated in Figure 2. Furthermore, our generation follows a structured storyline, similar to the human writing process, beginning with a brief outline and then drafting the complete story Yang et al. (2022b). This proposed task is complex and cannot be effectively solved by traditional frameworks or evaluated by existing benchmarks. A possible approach is to utilize powerful Multi-modal Large Language Models (MLLMs) Han et al. (2023); Sun et al. (2023) for zero-shot or few-shot generation. However, most MLLMs face challenges with time alignment and sequential video comprehension Huang et al. (2023). Moreover, generating narrations for longer videos becomes difficult for MLLMs due to the requirement of excessively long visual tokens. Another approach is to generate captions for video shots and use them as input for the LLMs Bhattacharya et al. (2023); Lin et al. (2023b). However, this relies heavily on the performance of the captioning model. Furthermore, current models are ineffective in generating coherent, appealing, and length-appropriate narrations. Additional training on specific instruction datasets is necessary.
To support the exploration of the proposed task, we create a benchmark dataset named E-SyncVidStory (E-commerce Synchronized Video Storytelling). Given the rapid growth of e-commerce platforms, many sellers have product information and visual materials, but may lack the time and resources to convert these into engaging advertisement videos. Therefore, they seek an automatic and efficient solution. Our dataset can provide value for both research and practical applications in this area. E-SyncVidStory contains 6.0K videos with 41.3K video clips in total. Each video is segmented as sequential visual scenes with synchronized Chinese narrations. As the primary objective of Ad stories is to engage with the audience and showcase products Bhattacharya et al. (2023), they might not always present a sequence of events as a classic story. Still, like classic ones, advertising stories should be coherent and follow a logical storyline, which is annotated in our dataset [^1] The storyline is expressed by categorized script labels. These labels are divided into 12 types, with their definitions provided in Appendix A.4.. All annotations are obtained through automatic preprocessing, GPT-based refinement, and manual checks.
@@ 3.1 Task Definition
Given a video $V$ composed of sequential video clips $\{v_{1},v_{2},...,v_{n}\}$ and the related knowledge composed of several knowledge items $\mathcal{K}=\{k_{1},k_{2},...,k_{m}\}$, synchronized video storytelling aims to incorporate the knowledge to generate a narration $s_{i}$ for each video clip $v_{i}$, and all narrations form a coherent story $\mathbf{S}=\{s_{1},s_{2},...,s_{n}\}$ for the video. The word count  [^2] If the time length of video clip $v_{i}$ is $t$ seconds, considering the normal speaking speed of 5 words per second, the appropriate word length range can be $[5t-3,5t+3]$ with an error margin of 3 words in either direction. of each $s_{i}$ should be within the appropriate range $r_{i}$. To enhance the coherence of the overall story, the generation is guided by a predefined or automatically predicted storyline, which is expressed using a sequence of script labels $\{l_{1},l_{2},...,l_{n}\}$. Each script label $l_{i}$, belonging to one of the 12 predefined types, summarizes the key point of the corresponding narration. See Appendix A.4 for script label details.
@@ 4.1 Visual Embedding
For long videos consisting of multiple video clips, when generating a narration for the current clip $v_{i}$, we only need a more concise message of previous video clips. Inspired by Song et al. (2023), we perform memory consolidation by merging the most similar visual tokens. Specifically, we maintain fixed-length tokens as visual memory. We iteratively repeat the merging process  [^4] Assume {$f_{i}$}${}_{i=1}^{k}$ are the visual features of k frames. In each merging process, we : (1) Evaluate the cosine similarity of each adjacent pair. (2) Find the pair ($f_{m},f_{m+1}$) with the highest similarity. (3) Merge $f_{m}$ and $f_{m+1}$ by averaging their features and remove $f_{m+1}$. until the memory tokens of previous video clips reach the fixed length.
@@
4.2 Prompting
The input prompt for VideoNarrator is displayed in Figure 6. For each video clip, the controllable signal is provided just before the narration, which can ensure controllable accuracy. In our instruction, we concisely explain the task requirements. As shown in Figure 6, these red words can make it easier for the LLM to target relevant information. In the underlined instruction, we emphasize that the model should primarily use the information provided in the prompt and avoid using unrelated knowledge stored in the LLM.
@@ 5.1 Experiment Settings
In our experiments, we use pre-trained models as visual feature extractors as described in the above Section. In experiments on E-SyncVidStory (advertising domain), we employ the Baichuan-7B model Yang et al. (2023) as the LLM model, known for its excellent performance in Chinese-related tasks. We construct our dataset as illustrated in Figure 6, resulting in 41k instruction samples. We apply 90% of them as the training set and the left as the testing set. Our proposed framework is trained for 15 epochs, using a learning rate of $1e^{-4}$ and a batch size of 32. The LoRA parameters are set to $r=64$ and $alpha=16$. The training of our 7B model took around $48$ hours on 4 A6000 GPUs.
In experiments on Video Storytelling dataset (general domain) Li et al. (2019), the LLM model is replaced by LLaMA-2 Touvron et al. (2023) for English generation. We make minor adjustments to fit the task setting of Synchronized Video Storytelling. Specifically, we apply the text length of the ground truth as the length requirement, and use the video topic as its relevant knowledge. The training prompt is similar to that presented in Figure 6, while the system role is changed to “You are an AI assistant that can understand videos”, the product information is replaced with the event topic, and the requirement for script labels has been removed.
@@
We compare our model with the following baselines. (1) Multi-model Pipieline. Specifically, we apply LLaVA Liu et al. (2023) to get textual summaries for video clips and prompting GPT-3.5 for zero-shot and few-shot generation. (2) End2end MLLMs, including zero-shot performance of Video-ChatGPT Maaz et al. (2023a), Video-LLaVA Lin et al. (2023a), and VTimeLLM Huang et al. (2023) with 7B LLMs. These models take the video as inputs and generate narrations in an end2end manner. (3) Fine-tuned End2end MLLMs. Specifically, we finetune the VTimeLLM model, which is effective for dense video comprehension and supports Chinese generation. More details about these baselines can be found in our Appendix A.1.
As for experiments on the Video Storytelling dataset Li et al. (2019), we compare our model with state-of-the-art methods: the retrieval-based GVMF Lu and Wu (2022) and the generative-based VerbalizeVideos Bhattacharya et al. (2023).
@@
Figure 7: Examples generated by our proposed VideoNarrator.
@@ 5.2 Automatic Evaluation Results
With Generated Storyline. As shown in Table 2, our model outperforms the baseline models, even when fine-tuned on our dataset. For detailed evaluations shown in Table 3, existing MLLMs demonstrate their ability to generate visually relevant and informative stories. However, they tend to focus on a limited amount of prior knowledge and may repeat it, resulting in a lower $\textrm{Info}_{\textrm{Diverse}}$ score. Furthermore, their fluency and ability to control the length of word sequences are poor. Based on the ablation studies, we find that by fine-tuning the LLM with our instruction data, our model can achieve the ability to control the word sequence length and comprehend video features, showing an increase in controllable accuracy and visual relevance. Applying visual compression and visual memory consolidation can improve the model’s ability to extract relevant knowledge and retain story coherence, resulting in better performance in knowledge relevance and fluency. Explicit storylines and relative positional embedding can encourage narrations to follow a more logical outline of persuasion, improving the performance in knowledge relevance.
@@
To validate the effectiveness of our proposed VideoNarrator, we conduct experiments on the Video Storytelling dataset Li et al. (2019), which covers a more general domain. As shown in Table 5, our model surpasses the state-of-the-art results. This showcases its effectiveness in synchronized video storytelling across a broader domain.
@@ 6 Conclusion
This paper introduces the new task of synchronized video storytelling, which aims to generate synchronous narrations for sequential video scenes. These narrations are guided by a structured storyline and can incorporate relevant knowledge, resulting in a coherent and informative story. The new task is more practical than existing tasks of dense video captioning and visual storytelling because the generated narrations can be directly combined with the visual scenes, creating an engaging video-format story. We collect a benchmark dataset called E-SyncVidStory and introduce an effective framework named VideoNarrator. Both automatic and human evaluations verify the effectiveness of our proposed framework.
@@
After refining the data using the powerful GPT-4, we only need to conduct further checking of the annotations. We have hired 5 workers to review and correct any remaining errors. Each video took approximately 2 minutes to complete, and the workers received reasonable payment for the local area.
Acknowledgements
We thank all reviewers for their insightful comments and suggestions. This work was partially supported by the National Natural Science Foundation of China (No. 62072462) and the Beijing Natural Science Foundation (No. L233008).
@@ A.2 Evaluation Metric for Knowledge Relevance
where $W(s_{i})$ represents all the words in sentence $s_{i}$, $T$ refers to the duration of the whole video. $f_{s_{i}},f_{k}$, and $f_{w}$ refer to the normalized embeddings of sentence $s_{i}$, knowledge point $k$, and segmented word $w$, respectively.